OpenStudy (anonymous):

hat does x in the solution equal in the following system of equations? –2x + y = 5 y = 2 – x

OpenStudy (anonymous):

x = -1

OpenStudy (anonymous):

Subtracting the second equation from the first,\[-2x+y-y=5-2+x\]\[-2x=3+x\]\[-3x=3\]\[x=-1\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

y = 2 – x , y+x=2 time(2)=2y+2x=4 2y+2x=4 y -2x =5 ADD-------------- 3y=9 y=9/3 y=3 Substitution y=3 find x –2x + y = 5 –2x + 3 = 5 -2x=5-3 x=2/-2 x=-1